OpenStudy (anonymous):

Graph the function f(x) = (x + 3)3 by hand and describe the end behavior.

OpenStudy (ash2326):

@Tati_Lee we have \[f(x)=(x+3)^3\] do you know the behavior of cubic functions?

OpenStudy (anonymous):

well i graphed it, i just need help describing the end behavior

OpenStudy (ash2326):

it's easy as \[x\to \infty, f(x) \to \infty\] and when \[x\to -\infty, f(x) \to -\infty\]

OpenStudy (ash2326):



OpenStudy (anonymous):

ohhh okay thanks

OpenStudy (ash2326):

did you understand?

OpenStudy (anonymous):

yess